# A Simple, Optimal and Efficient Algorithm for Online Exp-Concave Optimization

### v1

@@ C.1 Proof of Theorem˜4
First, we prove Eq.˜19. Without loss of generality, we consider $T\geq 3$. Then the high-probability excess risk bound of Eq.˜19 directly follows from Corollary 2 of Mehta [2017] by substituting LightONS’s regret in Theorem˜2.
@@
The first inequality is essentially Corollary 2 of Mehta [2017], and the second inequality follows from the fact that $\sqrt{ab}\leq\frac{a+b}{2}=O(a+b)$ for any positive terms $a$ and $b$. Recalling that $\textsc{Reg}_{T}=O(d\log T)=O\left(d\log\frac{d}{\epsilon}+d\log\log\frac{1}{\delta}\right)$, we have
$O\left(\frac{\textsc{Reg}_{T}}{T}\right)=O\left(\frac{d\left(\log\frac{d}{\epsilon}+\log\log\frac{1}{\delta}\right)}{\frac{d}{\epsilon}\log\frac{d}{\epsilon}\log\frac{1}{\delta}}\right)=O(\epsilon),$
and
@@
Next, we prove Eq.˜20 from Eq.˜19. By the definition of expectation, we have
$\mathop{\mathbb{E}}\left[F(\bar{\mathbf{x}}_{T^{\prime}})-\min_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})\right]$ $\leq O\left(\frac{\textsc{Reg}_{T^{\prime}}+\sqrt{\textsc{Reg}_{T^{\prime}}\cdot\log\Gamma_{1/T^{\prime}}}+\Gamma_{1/T^{\prime}}}{T^{\prime}}\right)+\frac{\max\limits_{\mathbf{y}\in\mathcal{X}}F(\mathbf{y})-\min\limits_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})}{T^{\prime}}$
$\leq O\left(\frac{\textsc{Reg}_{T^{\prime}}+\Gamma_{1/T^{\prime}}}{T^{\prime}}\right)+\frac{DG}{T^{\prime}}\leq O\left(\frac{\textsc{Reg}_{T^{\prime}}+\Gamma_{1/T^{\prime}}}{T^{\prime}}\right)$
@@
Finally, the total runtime follows from Theorem˜2 by substituting $T=\Theta\left(\frac{d}{\epsilon}\log\frac{d}{\epsilon}\log\frac{1}{\delta}\right)$. ∎
C.2 Proofs for ERM-based SXO Methods
In this part, we construct a scenario showing that ERM-based SXO methods cannot break the $\widetilde{O}(d^{3}/\epsilon)$ runtime barrier, even when equipped with state-of-the-art offline convex optimization solvers.
@@ C.2 Proofs for ERM-based SXO Methods
Under Assumption˜4 and $\mathcal{X}=\mathcal{B}(1)$, let the stochastic functions take the form $f(\mathbf{x};\xi)=\phi\left(\mathbf{w}(\xi)^{\mathsf{T}}\mathbf{x}\right)$, where $\phi:\mathbb{R}\to\mathbb{R}$ is a black-box function, and $\mathbf{w}:\Xi\to\mathbb{R}^{d}$ is a fixed mapping. by [Koren and Levy, 2015, Mehta, 2017], obtaining an $\epsilon$-optimal solution for SXO reduces to obtaining an $O(\epsilon)$-optimal solution to the offline ($\alpha$-exp-concave) objective
@@
CPM of Lee et al. [2015], Jiang et al. [2020] queries the gradient of $\widehat{F}$ for $\widetilde{O}(d)$ times, and each gradient query costs $O(dT)$ time due to the finite-sum structure of $\widehat{F}$ and the black-box nature of $\phi$. Therefore, the total runtime is $\widetilde{O}(d^{2}T+d^{3})=\widetilde{O}(d^{3}/\epsilon)$. ∎
Appendix D Proofs for Section˜5
In this section, we provide technical details on integrating LightONS into the applications in Section˜5 and their analysis.
D.1 Proofs for Gradient-Norm Adaptivity
The proof of Theorems˜6 and 5 closely follows that of their ONS counterparts with minor modifications.
Proof of Theorem˜5.
@@ D.1 Proofs for Gradient-Norm Adaptivity
Theorem˜2 implies the runtime with $\epsilon=d$ and the following regret bound:
@@
Then Corollary 5 of Orabona et al. [2012] converts the gradient-norm adaptive bound to the small-loss bound with respect to $L_{T}$ as defined Eq.˜22 and completes the proof. ∎
Proof of Theorem˜6.
Theorem 8 of Cutkosky and Orabona [2018] hinges on their Algorithm 7, Lemmas 16 and 17, apart from the coin-betting framework. To prove Theorem˜6, we show how LightONS adapts their analysis with minimal changes.
  • Modifications to their Algorithm 7. Since the decisions of their ONS are intermediate decisions to maximize the “wealth” in the coin-betting framework instead of true decisions to minimize regret, we can ignore the improper-to-proper conversion and replace their ONS with LightONS.Core. Specifically, their ONS runs on $\mathcal{X}=\mathcal{B}(1/2)$ while LightONS.Core runs on $\mathcal{Y}=\mathcal{B}(3/4)$ with the hysteresis coefficient $k=3/2$.
@@
  • Modifications to their Lemma 17. Relaxing the radius from $1/2$ to $3/4$ enlarges constants in the regret. Nonetheless, the numerical constants in their Lemma 17 are loose enough to accommodate our changes, greatly simplifying our analysis. Following their proof, let $\mathbf{z}_{t}$ denote the gradient that LightONS.Core receives at time $t$, Theorem˜2 and Eq.˜44 imply
@@
D.2 Proofs for Logistic Bandits

### v2

@@ D.1 Proof of Theorem˜3
First, we prove Eq.˜23a. Without loss of generality, we consider $T\geq 3$. Then the high-probability excess risk bound of Eq.˜23a directly follows from Corollary 2 of Mehta [2017] by substituting LightONS’s regret in Corollary˜1.
@@
The first inequality is essentially Corollary 2 of Mehta [2017], and the second inequality follows from the fact that $\sqrt{ab}\leq\frac{a+b}{2}=O(a+b)$ for any positive terms $a$ and $b$. Since $\textsc{Reg}_{T}=O(d\log T)=O\left(d\log\frac{d}{\epsilon}+d\log\log\frac{1}{\delta}\right)$, it suffices to show that
$O\left(\frac{\textsc{Reg}_{T}}{T}\right)$ $=O\left(\frac{d\left(\log\frac{d}{\epsilon}+\log\log\frac{1}{\delta}\right)}{\frac{d}{\epsilon}\log\frac{d}{\epsilon}\log\frac{1}{\delta}}\right)=O(\epsilon),\penalty 10000\ \penalty 10000\ \text{and}$
@@
Next, we prove Eq.˜24a from Eq.˜23a. By the definition of expectation, we have
$\mathop{\mathbb{E}}\left[F(\bar{\mathbf{x}}_{T^{\prime}})-\min_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})\right]$ $\leq\left(1-\frac{1}{T^{\prime}}\right)\cdot O\left(\frac{\textsc{Reg}_{T^{\prime}}+\sqrt{\textsc{Reg}_{T^{\prime}}\cdot\log\Gamma_{1/T^{\prime}}}+\Gamma_{1/T^{\prime}}}{T^{\prime}}\right)$
$\quad+\frac{1}{T^{\prime}}\left(\max_{\mathbf{y}\in\mathcal{X}}F(\mathbf{y})-\min_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})\right)$
$\leq O\left(\frac{\textsc{Reg}_{T^{\prime}}+\Gamma_{1/T^{\prime}}}{T^{\prime}}\right)+\frac{DG}{T^{\prime}}\leq O\left(\frac{\textsc{Reg}_{T^{\prime}}+\Gamma_{1/T^{\prime}}}{T^{\prime}}\right).$
@@
Finally, the total runtime in Eqs.˜23b and 24b follows from Theorem˜2. ∎
D.2 Proofs and Details for ERM-based SXO Methods
In this subsection, we provide two observations which support our conjecture in Section˜4.2 that ERM-based SXO methods fail to surpass the $\widetilde{O}(d^{3}/\epsilon)$ runtime barrier, even when these methods use more working memory. Moreover, we give an intuitive explanation of the runtime barrier $\widetilde{O}(d^{3}/\epsilon)$.
Fast matrix multiplication.
Linear regression with random design, a special case of SXO, reduces to solving a linear system with $d$ variables and $\widetilde{O}(d/\epsilon)$ equations. Although fast matrix multiplication accelerates this to $\widetilde{O}(d^{\omega}/\epsilon)$ runtime [Ibarra et al., 1982], this runtime does not plausibly extend beyond linear regression and reverts to $\widetilde{O}(d^{3}/\epsilon)$ in practice.
Cutting-plane methods.
The following proposition constructs an SXO instance where applying CPM correspondingly incurs $\widetilde{O}(d^{3}/\epsilon)$ runtime. We remark that, apart from the runtime, CPM further assumes well-roundedness of the domain, which may incur additional computational overhead.
@@ D.2 Proofs and Details for ERM-based SXO Methods
Under Assumption˜4 and that $\mathcal{X}=\mathcal{B}(1)$, let the stochastic functions take the form $f(\mathbf{x};\xi)=\phi\left(\mathbf{w}(\xi)^{\mathsf{T}}\mathbf{x}\right)$, where $\phi:\mathbb{R}\to\mathbb{R}$ is a black-box function, and $\mathbf{w}:\Xi\to\mathbb{R}^{d}$ is a fixed mapping. By [Koren and Levy, 2015, Mehta, 2017], obtaining an $\epsilon$-optimal solution for SXO reduces to obtaining an $O(\epsilon)$-optimal minimizer to the offline $\alpha$-exp-concave objective
@@
Best known CPM of Lee et al. [2015], Jiang et al. [2020] query the gradient of $\widehat{F}$ for $\widetilde{O}(d)$ times, and each gradient query costs $O(dT)$ time due to the finite-sum structure of $\widehat{F}$ and the black-box nature of $\phi$. Therefore, the total runtime is $\widetilde{O}(d^{2}T+d^{3})=\widetilde{O}(d^{3}/\epsilon)$. ∎
An intuitive explanation of the runtime $\widetilde{O}(d^{3}/\epsilon)$.
With the online-to-batch conversion, OGD’s regret $O(\sqrt{T})$ for OCO translates to a sample complexity of $O(1/\epsilon^{2})$ and a runtime of $\widetilde{O}(d/\epsilon^{2})$ for SXO. When the target excess risk $\epsilon$ is relatively large, i.e., $\epsilon=\widetilde{\Omega}(1/d^{2})$, the OGD-based SXO method may circumvent the $\widetilde{O}(d^{3}/\epsilon)$ runtime barrier. Thus the computational challenge of SXO mainly manifests when pursuing the (near) optimal sample complexity of $\widetilde{O}(d/\epsilon)$.
Lemma˜1 indicates that an exp-concave function can be regarded as a strongly convex function in the direction of its gradient. Comparing the sample complexity lower bound $\Omega(d/\epsilon)$ for SXO [Mahdavi et al., 2015] to the $\Omega(1/\epsilon)$ rate for stochastic strongly convex optimization [Hazan and Kale, 2011], the exp-concave offline objective $F$ can be viewed as the sum of $d$ strongly convex objectives on $d$ directions, one along each dimension corresponding to the eigenvectors of the Hessian-related matrix $A_{t}$ in ONS.
Our intuition is as follows, the difficulty of pursuing the (near) optimal sample complexity arises from exploiting the directional strong convexity, which necessitates estimating the covariance structure of the stochastic gradients. Cai et al. [2010] establish a minimax sample complexity for covariance matrix estimation under squared Frobenius norm, which coincides with the sample complexity lower bound for SXO. Specifically, for some family of covariance matrices $S\SS$, with $T$ samples i.i.d. sampled from a distribution with covariance $S\Sigma\in\SS$, the estimator $\widehat{\Sigma}$ satisfies
[…]
We refer readers to Theorem 1 of Cai et al. [2010] for a formal statement. Since computing the sample covariance matrix requires $O(d^{2}T)$ time, estimating the covariance structure of stochastic gradients plausibly incurs at least $O(d^{2}T)$ runtime, leading to the $\widetilde{O}(d^{3}/\epsilon)$ runtime barrier.
Appendix E Proofs for Section˜5
In this section, we provide algorithms and proofs for LightONS’s applications listed in Section˜5.
E.1 Proofs and Details for Gradient-Norm Adaptivity
$\textsc{Reg}_{T}(\mathbf{u})=O\left(d\log G_{T}\right),\quad\text{where}\penalty 10000\ G_{T}\triangleq\sum_{t=1}^{T}\left\|\nabla f_{t}(\mathbf{x}_{t})\right\|_{2}^{2}.$ (25)
In benign environments, $G_{T}$ can be $O(1)$, for example, when the loss is nearly fixed, decisions converge quickly, yielding regret far below worst-case bounds that scale with $T$. Yet, since $G_{T}\leq G^{2}T$, the gradient-norm adaptive bound safeguards the minimax-optimality in $T$. Prior work has leveraged Eq.˜25 to achieve such stronger types of adaptivity, including small-loss adaptivity for OXO with smoothness and comparator-norm adaptivity for unbounded OCO, as discussed below.
OXO with smoothness.
Orabona et al. [2012] demonstrate that ONS directly implies the desired regret bound in Eq.˜25 for OXO. Under smoothness assumptions, the gradient-norm adaptivity can further be improved to small-loss adaptivity [Srebro et al., 2010, Zhao et al., 2020], where the regret scales with the cumulative loss of the best comparator in hindsight, regardless of online decisions. This is formalized as follows:
$\textsc{Reg}_{T}=O\left(d\log L_{T}\right),\quad\text{where}\penalty 10000\ L_{T}\triangleq\min_{\mathbf{u}\in\mathcal{X}}\sum_{t=1}^{T}\left(f_{t}(\mathbf{u})-\min_{\mathbf{x}\in\mathcal{X}}f_{t}(\mathbf{x})\right),$ (26)
when the loss $f_{t}$ is $H$-smooth on $\mathcal{X}$ for any $t\in[T]$, i.e., $\left\|\nabla f(\mathbf{x})-\nabla f(\mathbf{y})\right\|_{2}\leq H\left\|\mathbf{x}-\mathbf{y}\right\|_{2}$ for any pair $(\mathbf{x},\mathbf{y})\in\mathcal{X}^{2}$.
Unbounded OCO.
Cutkosky and Orabona [2018] extend the gradient-norm adaptive OXO regret in Eq.˜25 to a comparator-norm adaptive OCO regret, i.e.,
$\textsc{Reg}_{T}(\mathbf{u})=\widetilde{O}\left(\left\|\mathbf{u}\right\|_{2}\sqrt{dG_{T}}\right),\quad\text{for any unbounded comparator}\penalty 10000\ \mathbf{u}\in\mathbb{R}^{d},$ (27)
without any prior knowledge of $\mathbf{u}$. Such guarantees for unbounded comparators are unachievable via classical OCO algorithms that explicitly depend on the domain diameter $D$, for example, OGD demands an explicit $D$ to obtain $O(DG\sqrt{T})$ regret [Zinkevich, 2003, Abernethy et al., 2008].
Nevertheless, both [Orabona et al., 2012, Cutkosky and Orabona, 2018] incur the worst-case $\widetilde{O}(d^{\omega}T)$ runtime bottleneck of ONS. Below, we show that LightONS can replace ONS in these algorithms to substantially improve efficiency without compromising statistical guarantees.
Improvements by LightONS.
LightONSrecovers ONS’s gradient-norm adaptive regret bound in Eq.˜25, thereby leading to small-loss adaptivity for OXO with smoothness, as well as comparator-norm adaptivity for unbounded OCO. We state the theoretical guarantees based on LightONS and omit those of the original ONS-based methods as they are identical except for the runtime.
Theorem 4 (LightONS’s improvement for OXO with smoothness).
Runtime $\leq O\left(\left(\texttt{EP}_{\mathcal{X}}+d^{2}\right)T+d^{\omega}\sqrt{T}\log T\right),$
where $L_{T}$ is defined in Eq.˜26, and $\gamma_{0}$ is defined in Corollary˜1.
Proof of Theorem˜4.
@@ E.1 Proofs and Details for Gradient-Norm Adaptivity
Theorem˜2 implies the runtime with $\epsilon=d$ and the regret of
@@
Then Corollary 5 of Orabona et al. [2012] converts the gradient-norm adaptive bound to the small-loss bound with respect to $L_{T}$ and completes the proof. ∎
Theorem 5 (LightONS’s improvement for unbounded OCO).
$\textsc{Reg}_{T}(\mathbf{u})$ $\leq\widetilde{O}\left(\sqrt{d\sum_{t=1}^{T}\left(\nabla f_{t}(\mathbf{x}_{t})^{\mathsf{T}}\mathbf{u}\right)^{2}}\right)\leq\widetilde{O}\left(\left\|\mathbf{u}\right\|_{2}\sqrt{dG_{T}}\right),$
Runtime $\leq O\left(d^{2}T+d^{\omega}\sqrt{T}\log T\right).$
Proof of Theorem˜5.
Theorem 8 of Cutkosky and Orabona [2018] hinges on their Algorithm 7, Lemmas 16 and 17, apart from the coin-betting framework. To prove Theorem˜5, we show how LightONS adapts their analysis with minimal changes.
  • Modifications to their Algorithm 7. Since the decisions of their ONS are intermediate decisions to maximize the “wealth” in the coin-betting framework instead of true decisions to minimize regret, we can ignore the improper-to-proper conversion and replace their ONS with LightONS.Core. Specifically, their ONS runs on $\mathcal{X}=\mathcal{B}(1/2)$ while LightONS.Core runs on $\mathcal{Y}=\mathcal{B}(3/4)$ with the deferral coefficient $k=3/2$.
@@
  • Modifications to their Lemma 17. Relaxing the radius from $1/2$ to $3/4$ enlarges constants in the regret. Nonetheless, the numerical constants in their Lemma 17 are loose enough to accommodate our changes, greatly simplifying our analysis. Following their proof, let $\mathbf{z}_{t}$ denote the gradient that LightONS.Core receives at time $t$, Theorem˜2 and Eq.˜28 imply
@@
Comparison with OQNS.
We remark that OQNS can hardly achieve full gradient-norm adaptivity, as the log-barrier regularization introduces an unavoidable $O(\log T)$ bias term independent of $G_{T}$. From the perspective of OMD, the regret can be decomposed into a stability term and a bias term. Adopted from Hazan [2016], Orabona [2019], the regret decomposition of a typical OMD algorithm is as follows:
While ONS and LightONS, as instances of OMD, can flexibly balance stability and bias,  [^3] For ONS and LightONS, the learning rate $\eta$ is replaced with the time-varying Hessian-related matrix’s inverse $A_{t}^{-1}$. OQNS sacrifices such flexibility for computational efficiency. OQNS employs a highly stable log-barrier, which suppresses the stability term but inflates the bias term in a manner resistant to small gradient norms. Specifically, the log-barrier contributes a term $\log(D^{2}-\left\|\mathbf{u}\right\|_{2}^{2})$ in the regret for any comparator $\mathbf{u}\in\mathcal{B}(D)$. As $\left\|\mathbf{u}\right\|_{2}$ approaches $D$, this term diverges to infinity. This issue is mitigated with the fixed-share trick [Mhammedi and Gatmiry, 2023], an illustration of which is as follows:
$\sum_{t=1}^{T}\left(f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{u})\right)\leq\underbrace{\sum_{t=1}^{T}\left(f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{v})\right)}_{\text{fixed-share regret}}+\underbrace{\sum_{t=1}^{T}\left(f_{t}(\mathbf{v})-f_{t}(\mathbf{u})\right)}_{\text{fixed-share margin}},\penalty 10000\ \penalty 10000\ \text{where}\penalty 10000\ \mathbf{v}=\left(1-\frac{1}{T}\right)\mathbf{u}.$
The fixed-share margin term $f_{t}(\mathbf{v})-f_{t}(\mathbf{u})\leq O(GD/T)$ due to boundedness and Lipschitzness, summing to $O(1)$ over $T$ rounds. Since $1-\left\|\mathbf{v}\right\|_{2}^{2}\geq\Omega(D^{2}/T)$, the fixed-share regret term contributes $\log(D^{2}-\left\|\mathbf{v}\right\|_{2}^{2})=O(\log T)$ to the regret, which prevents full gradient-norm adaptivity.
E.2 Proofs and Details for Logistic Bandits
In this subsection, we focus on logistic bandits as a representative instance of generalized linear bandits (GLB) to concretely illustrate the applicability of LightONS. GLB constitutes an important class of parametric stochastic bandits, where the expected loss depends on an unknown parameter through a known link function. This formulation introduces non-linearities into linear bandits, enhancing their expressivity while retaining tractable solutions.
Problem setting.
Logistic bandits can be interpreted as interactions between a learner and an adversary, which unfolds as follows: At each round $t\in[T]$, the learner selects an arm $\mathbf{x}_{t}$ from an arm set $\mathcal{X}_{t}\subseteq\mathcal{B}(1)^{K}$ with $K$ arms, and suffers a stochastic loss $y_{t}\in\{0,1\}$, where $\mathop{\mathbb{P}}\left(y_{t}=1\,|\,\mathbf{x}_{t}\right)=\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}^{\star})$. The true parameter $\mathbf{w}^{\star}\in\mathcal{W}=\mathcal{B}(D/2)$ is unknown, and $\sigma(z)=1/(1+\exp(-z))$ denotes the sigmoid function. The performance of the learner is measured by its pseudo-regret, quantifying the cumulative expected loss against the optimal arms in hindsight, which is defined as
A challenge in GLB lies in its dependence on $\kappa=\max_{\mathbf{x}\in\mathcal{X},\mathbf{w}\in\mathcal{W}}1/\sigma^{\prime}(\mathbf{x}^{\mathsf{T}}\mathbf{w})$, a problem-dependent constant that may grow expoenentially with $D$. Directly applying ONS to GLB yields a pseudo-regret of $\widetilde{O}(\kappa d\sqrt{T})$ [Jun et al., 2017], which becomes vacuous for $\kappa=\Omega(\sqrt{T})$.
Current results.
Recently, Zhang et al. [2025] propose the first one-pass GLB algorithm achieving $\kappa$-free-leading-term pseudo-regret with constant working memory and constant per-round time. Their pseudo-regret is $O(d\sqrt{T}\log T+\kappa(d\log T)^{2})$, which remains sublinear in $T$ for $\kappa=o(T)$.
Their key idea is to estimate the parameter $\mathbf{w}^{\star}$ with a “look-ahead” variant of ONS, whose analytical properties can remove the dependence on $\kappa$ in the pseudo-regret’s leading terms while retaining the constant-memory and constant-time efficiency of ONS. An illustration of their “look-ahead” ONS is provided as follows:
$\widetilde{H}_{t}=\epsilon I+\left(\sum_{i=1}^{t-1}\nabla^{2}\ell_{i}(\mathbf{w}_{i+1})\right)+\eta\nabla^{2}\ell_{t}(\mathbf{w}_{t}),\quad\widehat{\mathbf{w}}_{t+1}=\mathbf{w}_{t}-\frac{1}{\eta}\widetilde{H}_{t}^{-1}\nabla\ell_{t}(\mathbf{w}_{t}),\quad\mathbf{w}_{t+1}=\Pi^{\widetilde{H}_{t}}_{\mathcal{B}(D/2)}[\widehat{\mathbf{w}}_{t+1}],$ (30)
where $\ell_{t}(\mathbf{w})=-y_{t}\log\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w})-(1-y_{t})\log(1-\sigma(\mathbf{x}_{t}^{\mathsf{T}}\mathbf{w}))$ denotes the logistic loss for parameter estimation at round $t$.
Because of ONS, their algorithm incurs the worst-case $\widetilde{O}((d^{2}K+d^{\omega})T)$ runtime bottleneck, where $K$ is the number of arms. Below we show that plugging LightONS can improve the runtime while preserving the $\kappa$-free-leading-term pseudo-regret guarantee.
Improvements by LightONS.
We propose an analogous “look-ahead” variant of LightONS in Algorithm˜5, which replaces the ONS-based subroutine in Zhang et al. [2025]. Line 7 of Algorithm˜5 introduces the deferred-projection mechanism of LightONS.Core to replace the Mahalanobis projection in Eq.˜30.
Algorithm˜5 improves the runtime from $\widetilde{O}((d^{2}K+d^{\omega})T)$ to $\widetilde{O}(d^{2}KT+d^{\omega}\cdot\min\{\sqrt{\kappa dT},T\})$, while preserving the $\kappa$-free-leading-term pseudo-regret guarantee. When $\kappa=o(T)$, as required for sublinear pseudo-regret, the improved algorithm is asymptotically faster than the original algorithm. We state the theoretical guarantee of Algorithm˜5 in the theorem below.
Algorithm 5 “look-ahead” LightONS for [Zhang et al., 2025]
1: Initialize $H_{0}=\epsilon I$, $\mathbf{w}=\bm{0}$.
2:for $t=1,\dots,T$ do
5:  $\widetilde{H}_{t}=H_{t-1}+\eta\nabla^{2}\ell_{t}(\mathbf{w}_{t})$.
7:  $\mathbf{w}_{t+1}=\begin{cases}\widehat{\mathbf{w}}_{t+1}&\text{if}\penalty 10000\ \left\|\widehat{\mathbf{w}}_{t+1}\right\|_{2}\leq kD/2\\ \Pi^{\widetilde{H}_{t}}_{\mathcal{B}(D/2)}[\widehat{\mathbf{w}}_{t+1}]&\text{otherwise}\end{cases}$.
8:  $H_{t}=H_{t-1}+\nabla^{2}\ell_{t}(\mathbf{w}_{t+1})$.
9:end for
Theorem 6 (LightONS’s improvement for logistic bandits).